Analyzing Manufacturing Cost Accounts Fire Rock Company manufactures designer paddle boards in a wide variety of sizes and stWages Payable June 30 Wages incurred 121,300 Factory Overhead Factory overhead applied June 1 June 30 June 30 June 30 Balanceb. Factory overhead is applied to each job at a rate of 150% of direct labor cost. C. The June 1 Work in Process balance consJob No. Quantity June 1 Work in Process Direct Materials Direct Labor Factory Overhead Total Cost Unit Cost Units Sold Cost oa. Materials requisitions $ b. Work in process beginning balance $ c. Direct materials $ d. Direct labor $ e. Factory overhea
Answer #1

A)Material requisition = Direct material used to Jobs +Indirect material

           =136880+17100

           = 153980

b)Work in process beginning= 22500

c)Direct material = Material requisition - indirect material

              = 153980 -17100

                 = 136880

d)Direct labor = 97000

e)Factory overhead applied = Direct labor cost *% of overhead applied

                = 97000* 150%

                 = 145500

f)completed Jobs :333750

Job 201 202 204 205 Total
Beginning work in process 5000 17500 22500
Direct material 17920 41070 41100 18660 118750
Direct labor 13000 28000 30000 15000 86000
overhead applied 13000*150%= 19500 28000*150%= 42000

30000*150%= 45000

15000*150%= 22500 129000
Total cost 50420 111070 116100 56160 333750

G)cost of goods sold = 273399

Job No. Quantity Work in progress June 1 Direct material Direct labor overhead applied Total cost Unit cost[Total cost /quantity] Units sold cost of goods sold [unit cost*unit sold]
201 170 5000 17920 13000 19500 50420 296.58824 136 40336
202 430 17500 41070 28000 42000 111070 258.30233 344 88856
204 300 41100 30000 45000 116100 387 252 97524
205 160 18660 15000 22500 56160 351 133 46683
273399

H)

Indirect labor =wages payable -direct labor

          = 121300-97000

          = 24300